OpenStudy (anonymous):

evaluate each expression for each of these values:y=2, y=3, and y=5. y=2 y=3 y=5 a. y3 _______ _______ _______ b. 48 -y2 ________ _______ ________ c. 6y + 3y + 10 __________ ________ _________ d. (y+3)2 ____________ ___________ __________

OpenStudy (anonymous):

hope you can see it

OpenStudy (anonymous):

@Shay17

OpenStudy (anonymous):

Sorry I'm late hold on I have to write it down on paper first

OpenStudy (anonymous):

ok thats fine take your time

OpenStudy (anonymous):

For each expression substitute what y equals. In this case you have 4 expressions with 3 different y values so just substitute those y values in your expression Do you understand that ?

OpenStudy (anonymous):

yes i like the way you explain stuff

OpenStudy (anonymous):

So for a. you have y^3 and y equals 2 3 and 5 So that means 2^3 and 3^3 and 5^3

OpenStudy (anonymous):

Understand that now ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2^3 = 8 3^3 = 27 5^3 = 125

OpenStudy (anonymous):

Cool I'm glad you understand. This is the easy part about algebra there is a lot of stuff about algebra that is very hard. Lol

OpenStudy (anonymous):

lol thank you very much

OpenStudy (anonymous):

Let me know what you get for b c and d I wanna see your substitutions and the answers

OpenStudy (anonymous):

And no problem you are very welcome. It's my pleasure : )